# Generative AI and Firm Productivity:

### v2

@@ Abstract
We quantify the impact of Generative Artificial Intelligence (GenAI) on firm productivity through a series of large-scale randomized field experiments involving millions of users and products at a leading cross-border online retail platform. Over six months in 2023-2024, GenAI-based enhancements were integrated into seven consumer-facing business workflows. We find that GenAI adoption significantly increases sales, with treatment effects ranging from 0% to 16.3%, depending on GenAI’s marginal contribution relative to existing firm practices. Because inputs and prices were held constant across experimental arms, these gains map directly into total factor productivity improvements. Across the four GenAI applications with positive sales effects, the implied annual incremental value is approximately \$5 per consumer—an economically meaningful impact given the retailer’s scale and the early stage of GenAI adoption. The primary mechanism operates through higher conversion rates, consistent with GenAI reducing frictions in the marketplace and improving consumer experience. We also document substantial heterogeneity: smaller and newer sellers, as well as less experienced consumers, exhibit disproportionately larger gains. Our findings provide novel, large-scale causal evidence on the productivity effects of GenAI in online retail, highlighting both its immediate value and broader potential.
@@ 1 Introduction
The rapid diffusion of Generative Artificial Intelligence (GenAI) tools has sparked growing interest in their potential to reshape productivity across sectors of the economy. Recent academic research has provided compelling evidence of GenAI’s promise in various domains, including software development, customer support, education, and professional services [15, 54, 56, 26, 30, e.g.,]. Yet, despite rapid adoption, there is little empirical evidence of measurable gains in aggregate or firm-level revenue-based productivity attributable to GenAI [1]. Similarly, investors and industry practitioners have raised concerns about whether massive AI investments will translate into sustained business returns. [^1] See, for example, recent articles by Sequoia Capital: www.sequoiacap.com/article/ais-600b-question; and The Economist: www.economist.com/leaders/2025/09/11/what-if-the-3trn-ai-investment-boom-goes-wrong.
@@
Identifying the firm-level productivity impact of GenAI poses three empirical challenges. First, constraints in technical expertise and the need for complementary investments may delay implementation and the realization of observable gains, even when long-run potential exists [11]. [^2] For example, gains from technological innovations often take time to materialize [13]. Second, most existing applications of GenAI in firms remain at the pilot stage and focus on narrowly defined tasks—often at the worker level—making it difficult to detect productivity gains in aggregate firm-level data. [^3] For instance, recent studies have analyzed the effects of GenAI chatbots on workers’ performance [26, 55], and earnings [44]. See also: www.wsj.com/articles/companies-are-struggling-to-drive-a-return-on-ai-it-doesnt-have-to-be-that-way. This narrow scope overlooks a central insight from the economics literature: value creation stems not from isolated tasks, but from interdependent routines—or workflows—which play a central role in driving firm productivity [10, e.g.,]. Third, rigorous empirical analysis requires detailed revenue data and a setting that enables causal identification, both of which are rarely available.
@@
This paper provides large-scale, real-world experimental evidence on the causal impact of GenAI on productivity at both the firm and workflow levels, using data from one of the world’s largest cross-border online retail platforms. Over six months in 2023-2024, the platform integrated GenAI into seven consumer-facing business workflows—ranging from search query refinement to product description generation. In each workflow, GenAI augmented existing technologies with minimal or no displacement of labor and capital, ensuring that any observed changes in output reflect genuine productivity gains. [^4] In line with the classical Solow model [58], throughout the paper we interpret increases in output holding labor and capital constant as gains in total factor productivity. Each application was evaluated through randomized field experiments, with user groups ranging from tens of thousands to tens of millions. Leveraging granular consumer/product-level data, the experiments assess the short-term impact of GenAI on key performance outcomes such as sales (in dollar values) and conversion rates, allowing us to identify not only whether GenAI delivers measurable business outcomes and boosts productivity, but also where, how, and for whom those gains materialize.
@@
We document three main findings. First, most GenAI deployments generate economically significant gains, though the effects vary across workflows—from no detectable impact to increases of up to 16.3% in sales, with the largest improvements observed in customer service and search applications. Because output rose while labor and capital inputs remained constant, these improvements map directly into total factor productivity (TFP) gains of comparable magnitude. Aggregating across the four GenAI applications with positive sales effects, we estimate an annual incremental value of approximately \$5 per consumer. These impacts, observed both within and across workflows, are substantial given the retailer’s scale and the early stage of GenAI adoption. Second, the productivity gains arise from GenAI’s ability to reduce frictions in the online marketplace and increase purchase intention. Across workflows, we observe significantly higher conversion rates and no effect on average cart values, consistent with GenAI enhancing the consumer experience and inducing market expansion. Third, the effects are heterogeneous across seller and buyer segments: on both the demand and supply sides, less experienced and smaller buyers and sellers derive disproportionately larger gains from GenAI enhancements. Our setting represents markets across a diverse set of countries, languages, and cultures. This global scope makes the findings broadly generalizable to the online retail industry.
@@
A key distinction of our study relative to existing GenAI literature is its focus on revenue-based outcomes. In contrast to prior work emphasizing input-side efficiency gains—such as improvements in worker performance [15, 26, 56, e.g.,]—we show that firm-level adoption of GenAI can enhance productivity through demand-side value creation. In our setting, the estimated gains stem entirely from higher sales: across most experiments, workflow costs remained unchanged and GenAI deployment did not alter the factors of production. [^5] Section 3.3 provides details on the constant input structure of our experiments. Our estimates therefore represent a conservative lower bound on the potential returns on investments in GenAI. The observed improvements reflect enhanced consumer experience through GenAI-driven reductions in market frictions, as evidenced by higher conversion rates. These results are consistent with theories emphasizing that in consumer-facing sectors, quality improvements—such as product innovation, personalization, and relevance—can raise productivity without lowering input costs [61, 13, 25].
@@
Moreover, prior studies of GenAI have primarily examined its impact on individual-level productivity in narrowly defined tasks conducted in laboratory settings. Such isolated applications in diverse contexts make it difficult to compare results due to differences in implementation quality and organizational practices (see [18] for a review). Moreover, productivity dynamics in lab settings may not capture the complexities of real-world firm adoption, where technical, organizational, and market factors interact in intricate ways ([15]; [50]). In this study, we analyze the large-scale deployment of GenAI across multiple workflows—varying in business function and complexity—within a single firm. This setting allows us to assess GenAI’s impact on firm-level productivity, a dimension largely unexplored in prior work. By holding implementation and organizational factors constant, we can attribute productivity differences across workflows to how effectively GenAI augments each application’s baseline performance.
@@
We partnered with a world-leading cross-border e-commerce platform to identify and quantify the sources of GenAI-driven productivity gains in online retail. The platform enables consumers worldwide to buy directly from manufacturers at competitive prices. [^6] The platform connects hundreds of thousands of predominantly small-business sellers with hundreds of millions of active buyers across more than 100 countries and regions. The platform supports around 20 languages, providing localized services that facilitate global communication and accessibility. Between September 2023 and June 2024, the firm deployed GenAI solutions across seven consumer-facing business workflows: (1) Pre-sale Service Chatbot, (2) Search Query Refinement, (3) Product Description Generation, (4) Marketing Push Message Creation, (5) Google Advertising Title Optimization, (6) Chargeback Defense, and (7) Live Chat Translation. These correspond to three broader functional areas: (i) consumer and seller services (1, 6, 7); (ii) consumer–product matching (2, 3); and (iii) advertising and promotion (4, 5). Each workflow maps to a distinct stage of the customer journey, allowing us to assess GenAI’s productivity impact across a wide range of retail operations.
@@
Our setting involves multiple experiments applying GenAI to distinct business workflows. Several features enhance comparability across these experiments: all applications were developed and deployed by the same technical team—covering algorithm design, model fine-tuning, and online rollout—and operated within the same firm, under similar organizational and competitive conditions. Despite this common implementation environment, the effect of GenAI on workflow performance is likely to depend on the scope of its marginal contribution relative to baseline conditions, which differ across workflows. In each case, the control group reflects the firm’s standard practices prior to GenAI adoption. For instance, in the Pre-sale Service Chatbot, the control group received no customer support; in the Search Query Refinement, the search function in the control group relied on standard machine-learning-based search algorithms; and in most other workflows, the benchmark was human input.
@@
Within each workflow, GenAI deployment was evaluated through a large-scale randomized field experiment, comparing the GenAI-enhanced workflow to a baseline version used as the control. Notably, baseline workflows often included automation or human input but did not incorporate GenAI technologies. The treatment condition differed solely in the integration of GenAI, while prices and costs remained constant across conditions. Randomization occurred at the level of consumers—and in one case, products—with minimal overlap (less than 1%) across experiments. For five of the seven experiments, we obtained detailed consumer/product-level transaction data, including expenditure, conversions, and clicks. For analysis, we aggregate these data to the consumer level (and to the product level in one experiment) and leverage consumer, seller, and product characteristics to study treatment effect heterogeneity. Our primary outcome is sales value, measured by total consumer expenditure, which—given fixed prices and workflow costs in our setting—serves as a proxy for revenue-based productivity. We also examine conversion rates, a standard performance metric in online retail that captures consumer experience.
@@
Our results reveal that most GenAI deployments yield economically significant short-term productivity gains, though magnitudes differ across workflows. Among the five processes with detailed data, gains in sales range from no detectable effect in the advertising workflows to improvements of up to 16.3% in the Pre-sale Chatbot application, consistent with prior research on GenAI’s impact on individual tasks in lab settings (see, e.g., [56]). In the Search Query Refinement and Product Description workflows, the effects are smaller—generally 2-3%—yet still substantial for a platform of this scale and maturity. Back-of-the-envelope calculations based on the four deployments with detailed transaction data and positive effects—annualizing workflow-specific gains and assuming linear additivity—suggest that these GenAI applications generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects represent roughly 5.5–6% of the per-user revenue growth observed in global e-commerce between 2023 and 2024. We also document notable improvements in workflows without granular data: a 15% higher success rate in Chargeback Defense and a 5.2% increase in consumer satisfaction from Live Chat Translation.
@@
Taken together, these results show that GenAI generates sizable gains in targeted workflows and meaningful effects for a large and mature retailer, with further potential as adoption broadens and increasingly targets revenue-critical processes. For instance, while in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it had expanded to more than 40 applications and by 2025 to over 60. Meanwhile, API calls to proprietary GenAI tools increased twentyfold between 2024 and 2025, reflecting the rapid scaling of GenAI adoption across the platform. The long-run impact will ultimately depend on equilibrium forces—whether complementarities across workflows amplify these gains or industry-wide adoption offsets them through intensified competition.
We further shed light on the mechanisms underlying our results. We find that productivity improvements stem from enhanced consumer experience through GenAI-driven reductions in market frictions. Specifically, increased sales were strongly associated with higher conversion rates (extensive margins) and, where applicable, click-through rates, but not with higher average cart values (intensive margins). These conversion and engagement metrics increased by 1–22% across workflows, suggesting that GenAI applications contributed to productivity primarily through the creation of consumer value—most notably by improving the platform’s user experience and expanding the market rather than increasing spending per buyer. In particular, better pre-sale chatbots and richer product descriptions reduced information asymmetries; GenAI-refined queries lowered search frictions; and automated push messages closed gaps in content personalization. While platforms already alleviate such frictions [7], our results show that GenAI adoption can further mitigate them.
@@
Finally, we analyze heterogeneity in treatment effects across sellers, buyers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, larger gains should be expected among participants with lower baseline capabilities—namely, smaller and newer sellers, less experienced buyers with limited platform engagement, and products in the long tail or in less concentrated categories. Consistent with this view, we find that productivity gains are more pronounced for smaller and newer sellers, in line with recent evidence that GenAI particularly benefits users with lower baseline skills [15, 26]. On the demand side, the benefits are disproportionately greater for less experienced consumers. Overall, these patterns suggest that GenAI enhances the platform’s ability to reduce frictions between buyers and sellers, narrowing outcome gaps across participants. By contrast, the effects across product groups appear more context-dependent.
@@
The remainder of the paper is structured as follows. Section 2 reviews the relevant literature. Section 3 describes the context, theoretical framework, empirical framework, and data. Section 4 presents the aggregate results, followed by Section 5, which explores the heterogeneity of treatment effects. Section 6 concludes. The Appendix provides additional details and results.

### v3

@@ Abstract
We quantify the impact of Generative Artificial Intelligence (GenAI) on firm productivity through a series of large-scale randomized field experiments involving millions of users and products at a leading cross-border online retail platform. Over six months in 2023-2024, GenAI-based enhancements were integrated into seven consumer-facing business workflows. We find that GenAI adoption significantly increases sales, with treatment effects ranging from 0% to 16.3%, depending on GenAI’s marginal contribution relative to existing firm practices. Because inputs and prices were held constant across experimental arms, these gains map directly into total factor productivity improvements. Across the four GenAI applications with positive sales effects, the implied annual incremental value is approximately \$5 per consumer—an economically meaningful impact given the retailer’s scale and the early stage of GenAI adoption. The primary mechanism operates through higher conversion rates, consistent with GenAI reducing frictions and improving consumer experience. Importantly, these effects are not associated with worse post-purchase outcomes, as product return rates and customer ratings do not deteriorate. Finally, we document substantial demand-side heterogeneity, with larger gains for less experienced consumers. Our findings provide novel, large-scale causal evidence on the productivity effects of GenAI in online retail, highlighting both its immediate value and broader potential.
@@ 1 Introduction
The rapid diffusion of Generative Artificial Intelligence (GenAI) tools has sparked growing interest in their potential to reshape productivity across sectors of the economy. Recent academic research has provided compelling evidence of GenAI’s promise in various domains, including software development, customer support, education, and professional services [undefn, undefaaa, undefaac, undefy, undefac, e.g.,]. Yet, despite rapid adoption, there is little empirical evidence of measurable gains in aggregate or firm-level revenue-based productivity attributable to GenAI [undef]. Similarly, investors and industry practitioners have raised concerns about whether massive AI investments will translate into sustained business returns. [^1] See, for example, recent articles by Sequoia Capital: www.sequoiacap.com/article/ais-600b-question; and The Economist: www.economist.com/leaders/2025/09/11/what-if-the-3trn-ai-investment-boom-goes-wrong.
@@
Identifying the firm-level productivity impact of GenAI poses three empirical challenges. First, constraints in technical expertise and the need for complementary investments may delay implementation and the realization of observable gains, even when long-run potential exists [undefj]. [^2] For example, gains from technological innovations often take time to materialize [undefl]. Second, most existing applications of GenAI in firms remain at the pilot stage and focus on narrowly defined tasks, often at the worker level, making it difficult to detect productivity gains in aggregate firm-level data. [^3] For instance, recent studies have analyzed the effects of GenAI chatbots on workers’ performance [undefy, undefaab], and earnings [undefaq]. See also: www.wsj.com/articles/companies-are-struggling-to-drive-a-return-on-ai-it-doesnt-have-to-be-that-way. This task-level focus overlooks a central insight from the economics literature: productivity gains arise from changes in interdependent routines, or workflows, rather than from isolated tasks [undefi, e.g.,]. Third, rigorous empirical analysis requires detailed revenue data and a setting that enables causal identification, both of which are rarely available.
@@
This paper provides large-scale, real-world experimental evidence on the causal impact of GenAI on productivity at both the firm and workflow levels, using data from one of the world’s largest cross-border online retail platforms. Over a six-month period in 2023–2024, the platform integrated GenAI into seven consumer-facing business workflows, ranging from search query refinement to product description generation. In each workflow, GenAI augmented existing technologies with minimal or no displacement of labor and capital, ensuring that any observed changes in output reflect genuine productivity gains. [^4] In line with the classical Solow model [undefaae], throughout the paper we interpret increases in output holding labor and capital constant as gains in total factor productivity. Each application was evaluated through randomized field experiments, with sample sizes ranging from tens of thousands to tens of millions of users or products. Leveraging granular consumer- and product-level data, the experiments assess the short-term impact of GenAI on key performance outcomes such as sales (in dollar terms), conversion rates, and cart composition, allowing us to identify not only whether GenAI delivers measurable business value, but also where, how, and for whom these gains materialize.
@@
We document three main findings. First, most GenAI deployments generate economically significant gains, though the effects vary across workflows—from no detectable impact to increases of up to 16.3% in sales, with the largest improvements observed in customer service and search applications. Because output rose while labor and capital inputs remained constant, these improvements map directly into total factor productivity (TFP) gains of comparable magnitude. Aggregating across the four GenAI applications with positive sales effects, we estimate an annual incremental value of approximately \$5 per consumer. These impacts, observed both within and across workflows, are substantial given the retailer’s scale and the early stage of GenAI adoption. Second, productivity gains arise primarily from GenAI’s ability to reduce frictions along the customer journey and increase purchase probability. Across workflows, we observe significantly higher conversion rates but no effects on average cart values, product return rates, or customer ratings, consistent with GenAI improving the consumer experience and driving market expansion rather than altering spending intensity or misleading consumers. This mechanism is especially relevant in cross-country, multilingual online retail settings, where search and information frictions are common. Third, the effects of GenAI adoption are heterogeneous across market participants. On the demand side, less experienced consumers benefit disproportionately from GenAI-powered workflows. On the supply side, point estimates suggest larger gains for smaller and newer sellers, though these differences are estimated imprecisely and are often statistically insignificant.
@@
A key distinction of our study relative to the existing GenAI literature is its focus on revenue-based outcomes. In contrast to prior work emphasizing input-side efficiency gains, such as improvements in worker performance [undefn, undefy, undefaac, e.g.,], we show that firm-level adoption of GenAI can enhance productivity through demand-side value creation. In our setting, the estimated gains stem entirely from higher sales: across most experiments, workflow costs remain unchanged, and GenAI deployment does not alter the factors of production. [^5] Section 3.3 provides details on the constant input structure of our experiments. Our estimates therefore represent a conservative lower bound on the potential returns on investment in GenAI. The observed improvements reflect an enhanced consumer experience through GenAI-driven reductions in market frictions, as evidenced by higher conversion rates. These results are consistent with theories emphasizing that in consumer-facing sectors, quality improvements such as product innovation, personalization, and relevance can raise productivity without lowering input costs [undefaah, undefl, undefx].
@@
Moreover, prior studies of GenAI have primarily examined its impact on individual-level productivity in narrowly defined tasks, often conducted in laboratory or small-scale settings. Such isolated applications across diverse contexts make it difficult to compare results due to differences in implementation quality and organizational practices (see [undefq] for a review). In addition, productivity dynamics observed in lab settings may not capture the complexities of real-world firm adoption, where technical, organizational, and market factors interact in intricate ways [undefn, undefaw]. In this study, we analyze the large-scale deployment of GenAI across multiple workflows—varying in business function and complexity—within a single firm. This setting allows us to assess GenAI’s impact on firm-level productivity, a dimension that remains largely unexplored in prior work. By holding implementation and organizational factors constant, we can attribute productivity differences across workflows to how effectively GenAI augments each application’s baseline performance.
@@
We partnered with a world-leading cross-border e-commerce platform to identify and quantify the sources of GenAI-driven productivity gains in online retail. The platform enables consumers worldwide to purchase directly from manufacturers at competitive prices. [^6] The platform connects hundreds of thousands of predominantly small-business sellers with hundreds of millions of active buyers across more than 100 countries and regions. The platform supports around 20 languages, providing localized services that facilitate global communication and accessibility. Between September 2023 and June 2024, the firm deployed GenAI solutions across seven consumer-facing business workflows: (1) Pre-sale Service Chatbot, (2) Search Query Refinement, (3) Product Description Generation, (4) Marketing Push Message Creation, (5) Google Advertising Title Optimization, (6) Chargeback Defense, and (7) Live Chat Translation. These workflows span three broader functional areas: (i) consumer and seller services (1, 6, 7); (ii) consumer–product matching (2, 3); and (iii) advertising and promotion (4, 5). Each workflow corresponds to a distinct stage of the customer journey, allowing us to assess GenAI’s productivity impact across a wide range of retail operations.
@@
Our setting involves multiple experiments applying GenAI to distinct business workflows. Several features enhance comparability across these experiments: all applications were developed and deployed by the same technical team and operated within the same firm under similar organizational and competitive conditions. Despite this common implementation environment, the effect of GenAI on workflow performance is likely to depend on the scope of its marginal contribution relative to baseline conditions, which differ across workflows. In each case, the control group reflects the firm’s standard practices prior to GenAI adoption. For example, in the Pre-sale Service Chatbot workflow, the control group received no customer support; in Search Query Refinement, the control condition relied on standard machine-learning-based search algorithms; and in most other workflows, the benchmark was human input.
@@
Within each workflow, the GenAI deployment was evaluated through a large-scale randomized field experiment that compared the GenAI-enhanced workflow in the treatment group to a baseline version used in the control group. Notably, baseline workflows often included automation or human input but did not incorporate GenAI technologies. The treatment condition differed solely through the integration of GenAI, while prices and inputs remained constant across conditions. Randomization occurred at the level of consumers—and, in one case, products—with minimal overlap (less than 1%) across experiments, enabling causal identification of the effects of GenAI on consumer behavior. For five of the seven experiments, we obtained detailed consumer- or product-level transaction data, including expenditure, conversions, and clicks. For analysis, we aggregate these data to the consumer level (and to the product level in one experiment) and leverage consumer, seller, and product characteristics to study treatment-effect heterogeneity. Our primary outcome is sales value, measured by total consumer expenditure, which, given fixed prices and workflow costs in our setting, serves as a proxy for revenue-based productivity. We also examine conversion rates, a standard measure of consumer experience in online retail, as well as cart value, which captures changes in the behavior or preferences of consumers who make a purchase.
@@
Our results reveal that most GenAI deployments yield economically significant short-term productivity gains, though the magnitudes differ across workflows. Among the five processes with detailed data, gains in sales range from no detectable effect in the advertising workflows to improvements of up to 16.3% in the Pre-sale Service Chatbot, consistent with prior research on GenAI’s impact on individual tasks in lab settings (see, e.g., [undefaac]). In the Search Query Refinement and Product Description workflows, the effects are smaller, typically in the range of 2-3%, yet still substantial for a platform of this scale and maturity. Back-of-the-envelope calculations based on the four deployments with detailed transaction data and positive effects—annualizing workflow-specific gains and assuming linear additivity—suggest that these GenAI applications generate an annual incremental value of approximately \$4.6–\$5 per consumer. These effects correspond to roughly 5.5–6% of the per-user revenue growth observed in global e-commerce between 2023 and 2024. We also document notable improvements in workflows without granular data: a 15% increase in success rate for Chargeback Defense, and a 5.2% increase in consumer satisfaction from Live Chat Translation.
@@
Taken together, these results show that GenAI generates sizable gains in targeted workflows and meaningful effects for a large, mature retailer, with further potential as adoption broadens and targets revenue-critical processes. For example, while in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it had expanded to more than 40 applications and by 2025 to over 60. Meanwhile, API calls to proprietary GenAI tools increased twentyfold between 2024 and 2025, reflecting the rapid scaling of GenAI adoption across the platform. The long-run impact will ultimately depend on equilibrium forces, specifically whether complementarities across workflows amplify these gains or industry-wide adoption offsets them through intensified competition.
We further shed light on the mechanisms underlying our results. We find that productivity improvements stem primarily from enhanced consumer experience through GenAI-driven reductions in market frictions along the customer journey. Specifically, increases in sales are strongly associated with higher conversion rates—and, where applicable, improvements in intermediate engagement metrics such as click-through rates—but not with changes in spending intensity among buyers. Across workflows, conversion rates increase by 1–22%, while average cart values and purchase frequency, conditional on consumers making a purchase, remain unchanged. This pattern indicates that GenAI contributes to productivity mainly through market expansion, rather than through changes in spending behavior or preferences among existing buyers.
In addition, we find no evidence that these gains come at the expense of consumer welfare. Product return rates and customer ratings do not deteriorate following GenAI adoption, and in some cases, returns decline or ratings improve. Together, these findings suggest that GenAI-driven friction reduction improves decision quality and consumer experience rather than inducing deceptive or impulsive purchases. In particular, GenAI-powered pre-sale chatbots and richer product descriptions reduce search costs and information asymmetries; GenAI-refined queries lower search frictions and improve matching; and automated push messages enhance content personalization. While online platforms already mitigate such frictions [undeff], our results show that GenAI adoption can further reduce them along multiple stages of the customer journey.
@@
Finally, we analyze heterogeneity in treatment effects across buyers, sellers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, one would expect larger gains among participants with lower baseline capabilities, such as less experienced buyers with limited platform engagement, smaller and newer sellers, and products in the long tail or in less concentrated categories. Consistent with this view, we find robust and statistically significant heterogeneity on the demand side: less experienced consumers benefit disproportionately more from GenAI-powered workflows. On the supply side, point estimates suggest larger gains for smaller and newer sellers, though these differences are estimated imprecisely and are often statistically insignificant. Overall, these patterns suggest that GenAI enhances outcomes for participants with lower baseline capabilities. By contrast, effects across product groups are more context-dependent.
@@
The remainder of the paper is structured as follows. Section 2 reviews the relevant literature. Section 3 describes the context, theoretical and empirical frameworks, and data. Section 4 presents the main results, and Section 5 examines heterogeneity in treatment effects. Section 6 concludes. The Appendix provides additional details and results.